OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution

OpenStudy (anonymous):

man they really stretch to make these in to "word problems" don't they?

OpenStudy (anonymous):

haha yeah @satellite73

OpenStudy (anonymous):

\[2|x − 6| + 14 = 38\] subtract 14 get \[2|x − 6| = 24\] divide by 2 get \[|x-6|=12\]

OpenStudy (anonymous):

then solve \(x-6=12\) and \(x-6=-12\) separately, in one step by adding \(6\) to both sides have fun in lollipop world

OpenStudy (anonymous):

x = −6, x = 18 its this one @satellite73

OpenStudy (anonymous):

that is what i get too

OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution i dont think this one has a solution. @satellite73